Approximately 10% of all people are left-handed ("11 little-known facts," 2013). Consider a grouping of fifteen people.

e.) At least two are left-handed.

f.) At most three are left-handed.

g.) At least seven are left-handed.

h.) Seven of the last 15 U.S. Presidents were left-handed. Is this unusual? What does that tell you?
Answer #1

Here p= P(left handed) = 10% = 0.1

Given sample size (n ) =15

q = 1- p = 1 - 0.1 = 0.9

Let x be the no.of left handed persons

ere we have probability and fixed no.of independent trails

So, x can follow the binomial distribution